# Thermodynamic and observational implications of black holes in toroidal geometry

### v1

@@ III.3 Local Stability
$C_{P}(S_{\mathrm{HB}})=T_{\mathrm{HB}}\left(\frac{\partial S_{\mathrm{HB}}}{\partial T_{\mathrm{HB}}}\right)\Bigg{|}_{P,~Q}$ $=$ $\frac{2\sqrt{\pi}\left(2PS^{2}+3Q^{2}\right)}{3\sqrt{S}}\,.$ (19)
@@
$C_{P}(S_{\mathrm{Exp}})=T_{\mathrm{Exp}}\left(\frac{\partial S_{\mathrm{Exp}}}{\partial T_{\mathrm{Exp}}}\right)\Bigg{|}_{P,~Q}$ $=$ $\frac{2^{3/4}\sqrt{\pi}\left\{4P(S-1)+3Q^{2}\right\}}{3\sqrt[4]{S-1}}\,.$ (20)
@@
$C_{P}(S_{\mathrm{R}})=T_{\mathrm{R}}\left(\frac{\partial S_{\mathrm{R}}}{\partial T_{\mathrm{R}}}\right)\Bigg{|}_{P,~Q}$ $=$ $\frac{2\sqrt{\pi}\sqrt{\alpha}\left\{\frac{2P\left(e^{\alpha S}-1\right)^{2}}{\alpha^{2}}+3Q^{2}\right\}}{3\sqrt{e^{\alpha S}-1}}\,.$ (21)
@@
Figure 1: $C_{P}$ versus the Hawking-Bekenstein, exponential corrected and Rényi entropy $S$ by setting $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We obtain various curves for different entropy models to notice the impact of these entropies on $C_{V}$, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are demonstrated as red solid, blue dotted, and black dotted-dashed trajectories, respectively.
@@
Analyzing how specific heat varies with the size of a BH offers key insights into its stability under thermodynamic conditions and possible phase transitions. When the heat capacity is positive, the system tends to be stable and capable of equilibrating thermally with its surroundings; a negative value, on the other hand, points to instability, and a value of zero corresponds to a critical phase transition point. The heat capacity behavior in terms of different entropy frameworks is presented in Fig. 1 by setting $P=1,~\alpha=1$, and $Q=1$. Notably, we obtain different trajectories with respect to the entropy models that we employed in our analysis, such as for the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve), and Rènyi entropy (black dotted-dashed curve). We want to mention here that in the legends of all the thermodynamic figures, we present the Hawking-Bekenstein and exponential corrected entropies by HBE and Exp, respectively. It observes that heat capacity in terms of Hawking-Bekenstein and Rènyi entropies initially shows negative behavior for small values of entropy, and then they show positive behavior as the entropy rises, and also for small values, we obtain the zero point. For example, there is only one zero point or phase transition in the cases of the Hawking-Bekenstein (red solid curve) and Rènyi entropy (black dotted-dashed curve), which are at $0.7071$ and $0.5348$, respectively. However, in the case of exponential entropy (blue dotted curve), the heat capacity initially decreases with a positive behavior, and for small intervals of entropy, it becomes negative; however, after that, it again begins to grow with a positive behavior. In addition, we obtain two zero points at $1$ and $1.25$, indicating that in the presence of exponential entropy, we can obtain more zero points compared to other entropy frameworks. In BH thermodynamics, zero points act as pivotal points that distinguish between two unique thermal states; for example, they separate the unstable region, in which heat capacity shows negative behavior, from the stable region, in which heat capacity presents positive behavior. It is also often referred to as the restriction point in BH thermodynamic analysis, beyond which the system’s behavior becomes stable. It is essential to mention here that by changing the parameter’s value location of the zero point may also change; for example, in our case, if one chooses different values of Q,PQ,~P and $\alpha$ (in the case of Rènyi entropy), then the location of the zero point might vary too.
@@ III.4 PVP-V Diagram and Isotherm compressibility
$P(V_{\mathrm{HB}},~T_{\mathrm{HB}},~Q)$ $=$ $\frac{2\ 2^{2/3}\sqrt[6]{\pi}\left(\frac{3V}{4\sqrt{\pi}}+\sqrt{\pi}\right)}{3\sqrt[3]{3}V^{4/3}}\,.$ (22)
@@
$P(V_{\mathrm{Exp}},~T_{\mathrm{Exp}},~Q)$ $=$ $\frac{2^{3/4}\sqrt{\pi}\left\{4P(S-1)+3Q^{2}\right\}}{3\sqrt[4]{S-1}}\,.$ (23)
@@
$P(V_{\mathrm{R}},~T_{\mathrm{R}},~Q)$ $=$ $\frac{1}{9\ 2^{2/3}\alpha V^{2/3}+12\sqrt[3]{3\pi}}\bigg{(}\frac{4\sqrt[3]{2}(3\pi)^{2/3}\alpha Q^{2}}{V^{2/3}}+\frac{8\ 2^{2/3}\pi Q^{2}}{V^{4/3}}$ (24)
$+$ $\frac{6\ 2^{2/3}\sqrt{\alpha}T}{\sqrt{\alpha V^{2/3}}}\bigg{)}\,.$
@@
$K_{T}(\mathrm{HB})$ $=$ $\frac{9\sqrt[3]{6\pi}V^{4/3}}{16\pi Q^{2}+3TV}\,$, (26)
$K_{T}(\mathrm{Exp})$ $=$ $\frac{18(6\pi)^{2/3}V^{4/3}}{32\sqrt[3]{6}\pi^{4/3}Q^{2}-9T\left(V^{4/3}\right)^{5/4}}\,$, (27)
$K_{T}(\mathrm{R})$ $=$ $\left\{9\alpha\left(3\ 2^{2/3}\alpha V^{5/3}+4\sqrt[3]{3\pi}V\right)^{2}\right\}\bigg{\{}4\bigg{(}96\sqrt[3]{2}\pi\alpha^{2}Q^{2}V^{4/3}+32\ 2^{2/3}\sqrt[3]{3}\pi^{4/3}\alpha Q^{2}V^{2/3}$ (28)
$+$ $24(3\pi)^{2/3}\alpha^{3}Q^{2}V^{2}+\frac{3\sqrt[3]{2}\alpha^{3/2}TV^{2}\left(9\alpha V^{2/3}+2\sqrt[3]{6\pi}\right)}{\sqrt{\alpha V^{2/3}}}\bigg{)}\bigg{\}}^{-1}\,.$
@@ III.5 Helmholtz Free Energy
$\mathcal{F}(S_{\mathrm{HB}})$ $=$ $\frac{\sqrt{\pi}\left(9Q^{2}-2PS^{2}\right)}{3\sqrt{S}}\,.$ (30)
@@
$\mathcal{F}(S_{\mathrm{Exp}})$ $=$ $\frac{\sqrt{\pi}\left\{4P(S-4)(S-1)+3Q^{2}(5S-4)\right\}}{6\sqrt[4]{2}(S-1)^{5/4}}\,.$ (31)
@@
$\mathcal{F}(S_{\mathrm{R}})$ $=$ $\frac{2\left(e^{\alpha S}-1\right)\left\{2P\left(e^{\alpha S}-1\right)^{2}-\alpha^{2}Q^{2}\right\}}{2\alpha P\left(e^{\alpha S}-1\right)^{2}\left(3e^{\alpha S}-2\right)+\alpha^{3}Q^{2}\left(e^{\alpha S}+2\right)}\,.$ (32)
@@ III.6 Gibbs Free Energy
$\mathcal{G}(S_{\mathrm{HB}})$ $=$ $\frac{\sqrt{\pi}\left(2PS^{2}+9Q^{2}\right)}{3\sqrt{S}}\,.$ (34)
@@
$\mathcal{G}(S_{\mathrm{Exp}})$ $=$ $\frac{\sqrt{\pi}\left\{4P(S-1)(5S-8)+3Q^{2}(5S-4)\right\}}{6\sqrt[4]{2}(S-1)^{5/4}}\,.$ (35)
@@
$\mathcal{G}(S_{\mathrm{R}})$ $=$ $\frac{\sqrt{\pi}\left[3\alpha^{2}Q^{2}\left\{e^{\alpha S}(\alpha S+2)-2\right\}-2P\left(e^{\alpha S}-1\right)^{2}\left\{e^{\alpha S}(3\alpha S-4)+4\right\}\right]}{3\alpha^{3/2}\left(e^{\alpha S}-1\right)^{3/2}}\,.$ (36)
@@
In Fig. 4, we present the graphical behavior of GFE as a function of the Hawking-Bekenstein, exponential corrected entropy, and Rènyi entropies by putting Q=1,p=1Q=1,~p=1 and $\alpha=1$. One can notice that initially, GFE in terms of the Rènyi entropy (black dotted-dashed curve) decreases with positive behavior, but as the entropy increases, GFE continuously decreases with a negative trend, which indicates that in the Rènyi entropy case, the behavior of GFE is not stable. However, for both Hawking-Bekenstein (red solid curve) and exponential corrected (blue dotted curve) entropies, initially GFE decreases, but as entropy grows, GFE begins to rise with a positive trend, indicating the stability of GFE across all ranges of entropy.
@@ IV Thermodynamic Geometry in terms of Various entropy frameworks
This section presents different thermodynamic geometry formalisms (specifically, the Weinhold and Ruppiener metrics) based on various entropy corrections and examines how the divergence points align with the zero point (ZP) of the heat capacity Soroushfar:2020wch . The application of geometric frameworks through thermodynamic geometry has notably advanced our understanding of BH thermodynamic structures. In the context of thermal fluctuation theory, thermodynamic geometry provides an effective method for exploring the microscopic nature and phase transitions of BHs. We mention here that the curvature scalar, defined as an invariant over the thermodynamic space, provides essential insight into microscopic interactions and indicates critical behavior. This framework is frequently employed to study the nature of interactions within BH microstructures. The geothermodynamics methodology is utilized in this work to investigate the detailed connections between thermodynamic parameters and the corresponding geometric framework. We validate the phase transition for charged torus-like BHs by using different entropy frameworks. To begin with, we first define the basic framework of the thermodynamic geometry formalism. Therefore, we adopt the same process as described in Ref. Weinhold:1975xej , and the Weinhold metric framework is defined as
@@
Here, we mention that the Ricci curvature scalar obtained from the Weinhold metric formalism in terms of the Hawking-Bekenstein, exponential corrected and Rènyi entropies are zero which allows us to adopt another metric formalism that we employ in this study, is the Ruppeiner metric to investigate the coincidence of the divergence of the Ricci curvature scalar with the ZPs of heat capacity. In this metric formalism, we basically modify the Weinnhold metric given in Eq. (41) by multiplying it with the inverse temperature. The purpose of this scaling, from a physical perspective, is to ensure the consistency of the fluctuation theory underlying the Ruppeiner metric with temperature-dependent thermodynamic probability distributions. For example, the inclusion of inverse temperature enhances our understanding of curvature singularities as indicators of critical behavior or phase transitions in the context of BH thermodynamics. Therefore, the inverse temperature serves not just as a mathematical term but as a thermodynamically significant quantity that influences the structure of fluctuation geometry and reveals key aspects of stability and microscopic behavior. Thereby, one can define the Ruppeiner metric formalism given in Refs. Ruppeiner:2008kd ; Ruppeiner:1995zz ; Akbar:2011qw ; Soroushfar:2020wch , whose explicit form is given as
@@
$R^{\mathrm{Rup}}(S_{\mathrm{Exp}})=-\frac{-4P(S-1)-19Q^{2}}{8(S-1)\left\{Q^{2}-4P(S-1)\right\}}\,.$ (43)
@@
$R^{\mathrm{Rup}}(S_{\mathrm{R}})=-\frac{\alpha\left\{2P\left(5e^{\alpha S}-2\right)\left(e^{\alpha S}-1\right)^{2}+\alpha^{2}Q^{2}\left(15e^{\alpha S}+2\right)\right\}}{4\left(e^{\alpha S}-1\right)\left\{2P\left(e^{\alpha S}-1\right)^{2}-\alpha^{2}Q^{2}\right\}}\,.$ (44)
@@
One can obtain valuable insights regarding the possible phase transition and the microscopic interaction between the particles of the BHs by examining the Ricci curvature scalar. For example, a negative scalar curvature indicates strong attractive microscopic interactions, whereas a positive value suggests repulsive interactions between the particles. When the curvature is flat, it typically denotes either an ideal gas-like non-interacting system or a perfectly balanced interaction regime. The behavior of the curvature scalar obtained from the aforementioned formalism in various entropy frameworks by substituting Q=1,P=1Q=1,~P=1 and $\alpha=1$ is presented in Fig. 5. The trajectories that we plot in Fig. 5 are the Ricci scalar in terms of entropies such as the Hawking-Bekenstein (red solid curve), exponential corrected (blue dotted curve) and Rènyi entropies (black dotted-dashed curve). It observes that the ZPs of heat capacity also coincide with the divergence of the Ricci curvature scalar. In the Hawking-Bekenstein and Rènyi entropies cases, the Ricci curvature scalar $R^{\mathrm{Rup}}$ divergence coincides with the zero point of heat capacity at $0.7071$ and $0.5348$, respectively. Similarly, in the case of the exponential corrected entropy, the divergence of $R^{\mathrm{Rup}}$ coincides with both zero points of heat capacity at $1$ and $1.25$. One can also notice that the behavior of the Ricci curvature scalar is both negative and positive, which indicates that there is both attractive and repulsive interaction between the particles of the charged torus-like BHs.

### v2

@@ III.3 Local Stability
$C_{P}(S_{\mathrm{HB}})=T_{\mathrm{HB}}\left(\frac{\partial S_{\mathrm{HB}}}{\partial T_{\mathrm{HB}}}\right)\Bigg|_{P,~Q}$ $=$ $S\left(2-\frac{8\pi^{2}Q^{2}}{2PS^{2}+3\pi^{2}Q^{2}}\right)\,.$ (19)
@@
$C_{P}(S_{\mathrm{Exp}})=T_{\mathrm{Exp}}\left(\frac{\partial S_{\mathrm{Exp}}}{\partial T_{\mathrm{Exp}}}\right)\Bigg|_{P,~Q}$ $=$ $\frac{4(S-1)\left(\pi^{2}Q^{2}-4P(S-1)\right)}{4P(S-1)-5\pi^{2}Q^{2}}\,.$ (20)
@@
$C_{P}(S_{\mathrm{R}})=T_{\mathrm{R}}\left(\frac{\partial S_{\mathrm{R}}}{\partial T_{\mathrm{R}}}\right)\Bigg|_{P,~Q}$ $=$ $\frac{2\left(e^{\alpha S}-1\right)\left(2P\left(e^{\alpha S}-1\right)^{2}-\pi^{2}\alpha^{2}Q^{2}\right)}{2\alpha P\left(e^{\alpha S}-1\right)^{2}\left(3e^{\alpha S}-2\right)+\pi^{2}\alpha^{3}Q^{2}\left(e^{\alpha S}+2\right)}\,.$ (21)
@@
Figure 1: $C_{P}$ versus the Hawking-Bekenstein, exponential corrected and Rényi entropy $S$ by setting $P=1,~Q=1,~\alpha=1$ (in Rènyi entropy case). We obtain various curves for different entropy models to notice the impact of these entropies on $C_{P}$, such as the trajectories for the Hawking-Bekenstein, exponential corrected, and Rènyi entropies are demonstrated as red solid, blue dotted, and black dotted-dashed trajectories, respectively.
@@
Analyzing how specific heat varies with the size of a BH offers key insights into its stability under thermodynamic conditions and possible phase transitions. When the heat capacity is positive, the system tends to be stable and capable of equilibrating thermally with its surroundings; a negative value, on the other hand, points to instability, and a value of zero corresponds to a critical phase transition point. The heat capacity behavior in terms of different entropy frameworks is presented in Fig. 1 by setting $P=1,~\alpha=1$, and $Q=1$. Notably, we obtain different trajectories with respect to the entropy models that we employed in our analysis, such as for the Hawking-Bekenstein (red solid curve), exponential corrected entropy (blue dotted curve), and Rènyi entropy (black dotted-dashed curve). We want to mention here that in the legends of all the thermodynamic figures, we present the Hawking-Bekenstein and exponential corrected entropies by HBE and Exp, respectively. It observes that heat capacity in terms of Hawking-Bekenstein and Rènyi entropies initially shows negative behavior for small values of entropy, and then they show positive behavior as the entropy rises, and also for small values, we obtain ZP. For example, there is only one ZP or phase transition in the cases of the Hawking-Bekenstein (red solid curve) and Rènyi entropy (black dotted-dashed curve), which are at $2.2214$ and $1.1698$, respectively. However, in the case of exponential entropy (blue dotted curve), the heat capacity initially decreases with a positive behavior, and for small intervals of entropy, it becomes negative; however, after that, it again begins to grow with a positive behavior. In addition, we obtain two ZPs at $1$ and $3.4674$, indicating that in the presence of exponential entropy, we can obtain more ZPs compared to other entropy frameworks. In BH thermodynamics, ZPs act as pivotal points that distinguish between two unique thermal states; for example, they separate the unstable region, in which heat capacity shows negative behavior, from the stable region, in which heat capacity presents positive behavior. It is also often referred to as the restriction point in BH thermodynamic analysis, beyond which the system’s behavior becomes stable. It is essential to mention here that by changing the parameter’s value location of ZP may also change; for example, in our case, if one chooses different values of Q,PQ,~P and $\alpha$ (in the case of Rènyi entropy), then the location of ZP might vary too.
@@ III.4 PVP-V Diagram and Isotherm compressibility
$P(V_{\mathrm{HB}},~T_{\mathrm{HB}},~Q)$ $=$ $\frac{2\times\ 2^{2/3}\left(\pi Q^{2}+\frac{3\pi TV}{4}\right)}{3\sqrt[3]{3\pi}V^{4/3}}\,.$ (22)
@@
$P(V_{\mathrm{Exp}},~T_{\mathrm{Exp}},~Q)$ $=$ $\frac{8(6\pi)^{2/3}Q^{2}+9\sqrt[3]{6}\pi^{4/3}T\left(V^{4/3}\right)^{5/4}}{36V^{4/3}}\,.$ (23)
@@
$P(V_{\mathrm{R}},~T_{\mathrm{R}},~Q)$ $=$ $\frac{2(2\pi)^{2/3}}{3\sqrt[3]{3}V^{4/3}}\left(Q^{2}+\frac{3T\left(\alpha V^{2/3}\right)^{3/2}}{\alpha^{3/2}\left\{(6\pi)^{2/3}\alpha V^{2/3}+4\right\}}\right)\,.$ (24)
@@
$K_{T}(\mathrm{HB})$ $=$ $\frac{9\sqrt[3]{6}V^{4/3}}{\pi^{2/3}\left(16Q^{2}+3TV\right)}\,$, (26)
$K_{T}(\mathrm{Exp})$ $=$ $\frac{18\left(\frac{6}{\pi}\right)^{2/3}V^{4/3}}{32\sqrt[3]{6}Q^{2}-9\pi^{2/3}T\left(V^{4/3}\right)^{5/4}}\,$, (27)
$K_{T}(\mathrm{R})$ $=$ $\left[9\sqrt[3]{3}\sqrt{\alpha}V^{4/3}\left((6\pi)^{2/3}\alpha V^{2/3}+4\right)^{2}\right]\bigg[2(2\pi)^{2/3}\bigg\{8\sqrt{\alpha}Q^{2}\bigg(3\sqrt[3]{6}\pi^{4/3}\alpha^{2}V^{4/3}$ (28)
$+$ $4(6\pi)^{2/3}\alpha V^{2/3}+8\bigg)+\frac{3T\left(3(6\pi)^{2/3}\alpha V^{2/3}+4\right)\left(\alpha V^{2/3}\right)^{3/2}}{\alpha}\bigg\}\bigg]^{-1}\,.$
@@ III.5 Helmholtz Free Energy
$\mathcal{F}(S_{\mathrm{HB}})$ $=$ $\frac{3\pi Q^{2}}{\sqrt{S}}-\frac{2PS^{3/2}}{3\pi}\,.$ (30)
@@
$\mathcal{F}(S_{\mathrm{Exp}})$ $=$ $\frac{4P(S-4)(S-1)+3\pi^{2}Q^{2}(5S-4)}{6\sqrt[4]{2}\pi(S-1)^{5/4}}\,.$ (31)
@@
$\mathcal{F}(S_{\mathrm{R}})$ $=$ $\frac{\pi\sqrt{\alpha}Q^{2}\left(e^{\alpha S}(\alpha S+2)-2\right)}{\left(e^{\alpha S}-1\right)^{3/2}}-\frac{2P\sqrt{e^{\alpha S}-1}\left(e^{\alpha S}(3\alpha S-2)+2\right)}{3\pi\alpha^{3/2}}\,.$ (32)
@@ III.6 Gibbs Free Energy
$\mathcal{G}(S_{\mathrm{HB}})$ $=$ $\frac{6PS^{2}+17\pi^{2}Q^{2}}{4\pi^{2}Q^{2}S-8PS^{3}}\,.$ (34)
@@
$\mathcal{G}(S_{\mathrm{Exp}})$ $=$ $\frac{4P(S-1)(5S-8)+3\pi^{2}Q^{2}(5S-4)}{6\sqrt[4]{2}\pi(S-1)^{5/4}}\,.$ (35)
@@
$\mathcal{G}(S_{\mathrm{R}})$ $=$ $\frac{\pi\sqrt{\alpha}Q^{2}\left(e^{\alpha S}(\alpha S+2)-2\right)}{\left(e^{\alpha S}-1\right)^{3/2}}-\frac{2P\sqrt{e^{\alpha S}-1}\left(e^{\alpha S}(3\alpha S-4)+4\right)}{3\pi\alpha^{3/2}}\,.$ (36)
@@
In Fig. 4, we present the graphical behavior of GFE as a function of the Hawking-Bekenstein, exponential corrected entropy, and Rènyi entropies by putting Q=1,p=1Q=1,~p=1 and $\alpha=1$. One can notice that initially, GFE in terms of the Rènyi entropy (black dotted-dashed curve) decreases with positive behavior, but as the entropy increases, GFE continuously decreases with a negative trend, which indicates that in the Rènyi entropy case, the behavior of GFE is not stable. However, for both the Hawking-Bekenstein (red solid curve) and the exponential-corrected (blue dotted curve) entropies, GFE initially decreases, but as entropy increases, it begins to rise, indicating its stability across all entropy ranges.
Furthermore, the comparative analysis of all the above-mentioned thermodynamic quantities reveals that the exponential corrected entropy offers a more intricate phase structure, characterized by two specific heat ZPs absent in the Hawking-Bekenstein and Rènyi cases. This distinction reveals a more profound dependence of the microstructure on the horizon’s topology, which we explore geometrically in the following section by employing thermodynamic curvature.
@@ IV Thermodynamic Geometry in terms of Various entropy frameworks
This section presents different thermodynamic geometry formalisms (specifically, the Weinhold and Ruppiener metrics) based on various entropy corrections and examines how the divergence points align with ZP of the heat capacity Soroushfar:2020wch . The application of geometric frameworks through thermodynamic geometry has notably advanced our understanding of BH thermodynamic structures. In the context of thermal fluctuation theory, thermodynamic geometry provides an effective method for exploring the microscopic nature and phase transitions of BHs. We mention here that the curvature scalar, defined as an invariant over the thermodynamic space, provides essential insight into microscopic interactions and indicates critical behavior. This framework is frequently employed to study the nature of interactions within BH microstructures. The geothermodynamics methodology is utilized in this work to investigate the detailed connections between thermodynamic parameters and the corresponding geometric framework. We validate the phase transition for charged torus-like BHs by using different entropy frameworks. To begin with, we first define the basic framework of the thermodynamic geometry formalism. Therefore, we adopt the same process as described in Ref. Weinhold:1975xej , and the Weinhold metric framework is defined as
@@
Here, we mention that the Ricci curvature scalar obtained from the Weinhold metric formalism in terms of the Hawking-Bekenstein, exponential corrected and Rènyi entropies are zero (this signifies that the equilibrium manifold is flat in this metric formulation) which allows us to adopt another metric formalism that we employ in this study, is the Ruppeiner metric to investigate the coincidence of the divergence of the Ricci curvature scalar with the ZPs of heat capacity. In this metric formalism, we basically modify the Weinnhold metric given in Eq. (41) by multiplying it with the inverse temperature. The purpose of this scaling, from a physical perspective, is to ensure the consistency of the fluctuation theory underlying the Ruppeiner metric with temperature-dependent thermodynamic probability distributions. For example, the inclusion of inverse temperature enhances our understanding of curvature singularities as indicators of critical behavior or phase transitions in the context of BH thermodynamics. Therefore, the inverse temperature serves not just as a mathematical term but as a thermodynamically significant quantity that influences the structure of fluctuation geometry and reveals key aspects of stability and microscopic behavior. Thereby, one can define the Ruppeiner metric formalism given in Refs. Ruppeiner:2008kd ; Ruppeiner:1995zz ; Akbar:2011qw ; Soroushfar:2020wch , whose explicit form is given as
@@
$R^{\mathrm{Rup}}(S_{\mathrm{Exp}})=-\frac{-4P(S-1)-19\pi^{2}Q^{2}}{8(S-1)\left(\pi^{2}Q^{2}-4P(S-1)\right)}\,.$ (43)
@@
$R^{\mathrm{Rup}}(S_{\mathrm{R}})=-\frac{\alpha\left\{2P\left(5e^{\alpha S}-2\right)\left(e^{\alpha S}-1\right)^{2}+\pi^{2}\alpha^{2}Q^{2}\left(15e^{\alpha S}+2\right)\right\}}{4\left(e^{\alpha S}-1\right)\left\{2P\left(e^{\alpha S}-1\right)^{2}-\pi^{2}\alpha^{2}Q^{2}\right\}}\,.$ (44)
@@
One can obtain valuable insights regarding the possible phase transition and the microscopic interaction between the particles of the BHs by examining the Ricci curvature scalar. For example, a negative scalar curvature indicates strong attractive microscopic interactions, whereas a positive value suggests repulsive interactions between the particles. When the curvature is flat, it typically denotes either an ideal gas-like non-interacting system or a perfectly balanced interaction regime. The behavior of the curvature scalar obtained from the aforementioned formalism in various entropy frameworks by substituting Q=1,P=1Q=1,~P=1 and $\alpha=1$ is presented in Fig. 5. The trajectories that we plot in Fig. 5 are the Ricci scalar in terms of entropies such as the Hawking-Bekenstein (red solid curve), exponential corrected (blue dotted curve), and Rènyi entropies (black dotted-dashed curve). It observes that the ZPs of heat capacity also coincide with the divergence of the Ricci curvature scalar. In the Hawking-Bekenstein and Rènyi entropies cases, the Ricci curvature scalar $R^{\mathrm{Rup}}$ divergence coincides with ZP of heat capacity at $2.2214$ and $1.1698$, respectively. Similarly, in the case of the exponential corrected entropy, the divergence of $R^{\mathrm{Rup}}$ coincides with both ZPs of heat capacity at $1$ and $3.4674$.
One can also notice that the behavior of the Ricci curvature scalar is both negative and positive, which indicates that there is both attractive and repulsive interaction between the particles of the charged torus-like BHs that can be related to the non-trivial toroidal topology. Owing to the dual periodic identification of the $S^{*}\times S^{*}$ horizon, the fluctuation manifold gains an additional curvature freedom, enabling $R^{\mathrm{Rup}}$ to be highly sensitive to small changes in entropy and providing a dual transitional behavior. In a broader context, the singularities in $R^{\mathrm{Rup}}$ correspond to topological defects within the thermodynamic state manifold, where their numbers of divergence (or ZP) corresponds to the phase transition order; in our case (toroidal BH), the existance of two such divergence in $R^{\mathrm{Rup}}$ signifies a modification in the thermodynamic manifold’s topological charge, which also agrees with the recent studies on thermodynamic topology Wei:2020poh ; Bhattacharya:2019qxe ; Wei:2021vdx ; Jeon:2024yey . The Alignment of divergence points of $R^{\mathrm{Rup}}$ with the ZP of heat capacity confirms that geothermodynamic formalisms serve as a consistent and independent tool for diagnosing phase transitions.